OpenStudy (anonymous):

What is the x–intercept of 2x + 5y = –10?

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x 2x + 5y = -10 2x + 5*0 = -10 2x + 0 = -10 I'll let you finish

OpenStudy (anonymous):

(-5,0) thanks

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Yay thanks so much!!

OpenStudy (anonymous):

simply let y=0 then you will get 2x=-10 then u divide both sides by 2 u will get x= -5 ......hope is clear...

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Oh and thanks msamido